# MMSR: Symbolic Regression is a Multimodal Task

### v4

@@ Abstract
Mathematical formulas are the crystallization of human wisdom in exploring the laws of nature for thousands of years. Describing the complex laws of nature with a concise mathematical formula is a constant pursuit of scientists and a great challenge for artificial intelligence. This field is called symbolic regression. Symbolic regression was originally formulated as a combinatorial optimization problem, and GP and reinforcement learning algorithms were used to solve it. However, GP is sensitive to hyperparameters, and these two types of algorithms are inefficient. To solve this problem, researchers treat the mapping from data to expressions as a translation problem. And the corresponding large-scale pre-trained model is introduced. However, the data and expression skeletons do not have very clear word correspondences as the two languages do. Instead, they are more like two modalities (e.g., image and text). Therefore, in this paper, we proposed MMSR. The SR problem is solved as a pure multimodal problem, and contrastive learning is also introduced in the training process for modal alignment to facilitate later modal feature fusion. It is worth noting that to better promote the modal feature fusion, we adopt the strategy of training contrastive learning loss and other losses at the same time, which only needs one-step training, instead of training contrastive learning loss first and then training other losses. Because our experiments prove training together can make the feature extraction module and feature fusion module wearing-in better. Experimental results show that compared with multiple large-scale pre-training baselines, MMSR achieves the most advanced results on multiple mainstream datasets including SRBench.
Machine Learning, ICML
Equal contributionJingyi Liu
Figure 1: From the radar chart, we can see that the comprehensive performance of our algorithm achieves the state of the art on each data set.
@@ 1 Introduction
Symbolic Regression (SR) aims to identify the explicit mathematical expression underlying the observed data. Formally, for a given dataset $D=\{X,y\}$, SR seeks a function $f$ that satisfies $y=f(X)$, where $X\in\mathbb{R}^{n\times d}$, $y\in\mathbb{R}^{n}$, $d,n$ are the dimension of variable and number of data points, respectively. $f$ is composed of several basic primitive operators such as +,,×,÷+,-$\times$,$\div$+ , - , × , ÷, etc. SR has been applied in many fields, for example, material sciences(Wang et al., 2019), physical law exploration(Udrescu & Tegmark, 2020)(Udrescu et al., 2020)(Schmidt & Lipson, 2009), etc.
@@
Genetic Programming Symbolic Regression (GPSR) has dominated in the past. GPSR uses the pre-order traversal of expression and acts as genes, and genetic operations such as mutation and crossover are used to produce descendants. Constants are randomly generated or optimized with the PSO algorithm. Although they have shown distinguished performances in several public benchmarks, they are sensitive to hyperparameters (Petersen, 2019) and the complexity of the learned expression tends to be high with the epoch increased.
@@
With the huge success gained by deep learning in multiple domains (Alec et al., 2018), some works apply the deep learning technique in SR. EQL (Kim et al., 2021a) first substitutes the activation function of a fully connected feedforward neural network and refines the expression with Lasso. However, EQL is hard to tackle the division operation because it would generate exaggerated gradients in backpropagation. Therefore, DSR (Petersen, 2019) utilizes an RNN to emit the skeleton of expressions, constants are represented with a constant placeholder and optimized later by SGD. The risk-seeking policy gradients are used to update the RNN. To step further, DSO (Mundhenk et al., 2021) incorporates genetic programming into the search: RNN provides a better population for genetic programming algorithms. By doing so, DSO achieves remarkable results in SRBench (La Cava et al., 2021), Since the efficiency of reinforcement learning is lower than that of numerical optimization, MetaSymNet(Li et al., 2023) skillfully uses numerical optimization problem to solve the combinatorial optimization problem of symbolic regression.
@@
The above-stated methods belong to the end-to-end model which trains a new model for each problem. This type of training strategy makes their search time long. Plus, the model can’t make full use of the search experiences from other searches. Therefore, the large-scale pre-trained model is proposed to improve the search speed. Normally, the large-scale pre-trained models (Valipour et al., 2021)(Luca et al., 2021)(Vastl et al., 2022) are trained with enormous self-generated data which contains the pairwise of sampled data points and corresponding mathematical expressions. The neural network has an encoder-decoder structure, the encoder maps the input data points into data features and the decoder uses the feature to generate the pre-order traversal of expressions. The Beam Search (Kumar et al., 2013) technique is applied to produce multiple candidate skeletons. Then, Broyden-Fletcher-Goldfarb-Shanno (BFGS) (Liu & Nocedal, 1989) is used to optimize the constants of candidate skeletons. Finally, the best-performed expression is selected as the final result.
@@
The large-scale pre-trained model gained fast search speed because once the model finished training, we only needed a single forward pass to get the expressions. However, they ignore the different data formats of the input and output of SR. That is, previous methods treat the SR as a machine translation task: translate the data points into corresponding expressions. However, it is clear that the input and output of SR are very different from the two languages in translation, and they do not have obvious word correspondence like languages. They are more like two different modalities (like image and text). Therefore, it is perhaps more reasonable to think of SR as a multimodal task.
@@
To tackle the problems of two different modalities, we use a modal alignment module in our model to make the encoded data feature close to the skeleton feature. We name our model as MMSR. Specifically, we use the SetTransformer (Lee et al., 2019) as a data encoder for it presents permutation invariance. And, decodes the skeleton of expression with a causal masking transformer decoder. Unlike standard decoder transformers, MMSR omits cross-attention in the first half of the decoder layers to encode unimodal skeleton representations, and cascades the rest of the decoder layers, cross-attending to the data encoder for multimodal data-skeleton representations. We use contrastive learning (Chuang et al., 2020) to align the modal features, so that the matching data features and skeleton features are as close as possible in the feature space, and the non-matching features are as far away from each other as possible. Additionally, we train the model with two extra losses. First, the cross-entropy loss is used for predicting the token of expression skeletons. Second, one of the limitations of pre-train models is that they focus on the supervised pretraining goals borrowed from symbol generation, i.e., they are trained solely with the token-level cross-entropy (CE) loss, which can result in equations that may exhibit high token-level similarities but are suboptimal with respect to equation-specific objectives such as fitting accuracy(Shojaee et al., 2024). To solve this problem, we directly predict the numerical value of the constants and use MSE to measure the loss between the predicted constants and the true constants in the skeletons. We summarize our contributions as follows:
  • We try to treat the symbolic regression task as a purely multimodal task according to the characteristics of the input and output of symbolic regression. And, a single end-to-end training session suffices to directly generate satisfactory outcomes.
  • We introduce a modality alignment method so that the matched data features and skeleton features in the feature space can be as close as possible, and the mismatched features can be as far away as possible. To facilitate the later feature fusion. In particular, our contrastive loss and cross-entropy loss are both done in one step instead of separately trained. Experiments show that this can better promote the fusion of features.
  • We conducted plenty of experiments and the results showed that the proposed model achieved state-of-the-art performances against the benchmarks.
@@ 2.1 Multimodal
Recently, models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) performed pre-train on noisy images and text pairs from the web using contrastive loss. Contrastive loss is one of the most effective feature learning methods(He et al., 2020)(Chen et al., 2020)(Li et al., 2020b)(Li et al., 2020a). They achieve remarkable performance on image-text retrieval tasks but lack the ability to model more complex interactions between image and text for other V+L tasks (Kim et al., 2021b). (e.g. VQA(Antol et al., 2015)). Subsequent investigations (Wang et al., 2021) (Wang et al., 2022) (Piergiovanni et al., 2022) have introduced encoder-decoder frameworks trained utilizing generative loss functions, which demonstrate robust performance across vision-language benchmarks. Concurrently, the visual encoders within these models maintain competitive accuracy in image classification tasks. Studies (Singh et al., 2022) (Li et al., 2021) (Li et al., 2022a)(Chen et al., 2023)(Liu et al., 2024) have investigated the unification of image and text representations, which require multiple pretraining stages of unimodal and multimodal modules to achieve high-performance levels. For example, ALBEF (Li et al., 2021) employs a dual-encoder architecture that integrates contrastive loss with Masked Language Modeling (MLM) to enhance learning efficiency. CoCa(Yu et al., ) focuses on training an image-text foundation model from scratch in a single pretraining stage to unify these approaches. This is simpler and more efficient to train.
@@ 2.2 Deep Symbolic Regression
Before the explosion of deep symbolic regression methods, Genetic Programming Symbolic Regression (GPSR) (Arnaldo et al., 2014)(McConaghy, 2011)(Nguyen et al., 2017)(Zhang et al., 2021) was the main strategy for solving symbolic regression tasks. GPSR encodes the pre-order traversal of mathematical expression as the gene type, by randomly initializing a population of individuals, GPSR generates the candidate set of expressions. Then, it uses a fitness function to evaluate the fitting results, and genetic operations such as mutation and crossover are applied to produce new generations. The algorithm iterates until reaches the expected fitness value or the pre-defined maximum iteration epoch. Although GPSR could perform well in public benchmarks, it is reported to be sensitive to hyperparameters (Petersen, 2019). Furthermore, the GP-based method lacks an effective way to determine the constant value in the expression.
Figure 2: Our model’s framework.
@@
With the huge success achieved by deep learning in various domains, EQL (Equation Learner) revised the activation functions of fully connected neural networks to primitives (sin, cos, etc.) and obtained the skeleton and constants through Lasso (Tibshirani, 1996) technique. However, the nontraditional activation functions cause troubles to the backpropagation process: the division operator would cause infinite gradients thus making the weights updating hard to continue. Moreover, the primitive sin, cos are periodical, and log, sqrt are not differentiable everywhere. DSR (Petersen, 2019) takes another learning strategy: it learns the skeleton first and optimizes the constants later. DSR uses an RNN to emit the probability distribution of the pre-order traversal, and autoregressively sample expressions according to the distribution. For gradient information lost in the sampling process, reinforcement learning is applied to update the RNN. Specifically, the risk-seeking policy gradients are used. Furthermore, to utilize the strong ability to generate new performed well expressions of GPSR, DSO (Mundhenk et al., 2021) incorporating the DSR and GPSR method to achieve state-of-the-art performance in SRBench. The RNN could provide a better population for GPSR to produce new descendants.
@@
DSO is a type of learning-from-scratch method, it needs to train a new model for every problem. This training strategy comes with a huge search time cost. Thus, many learning-with-experience methods are emerged. They usually employ a data encoder to receive the data points as inputs and output the expression encoding as the label to train a large-scale model. When a new problem needs to be handled, the learning-with-experience method just implements a forward propagation to obtain the expressions thus being time-effective. SymbolicGPT (Valipour et al., 2021) treats the problem as a natural language problem, it uses symbolic-level encoding as the label of a set of data points. For example, the expression $2sin(x+y)$ is represented as text $[^{\prime}2^{\prime},^{\prime}s^{\prime},^{\prime}i^{\prime},^{\prime}n^{\prime},^{\prime}(^{\prime},^{\prime}x^{\prime},^{\prime}+^{\prime},^{\prime}y^{\prime},^{\prime})^{\prime}]$. This type of encoding forces the model to learn the syntax and semantics of the primitives thus increasing the learning difficulty. To free the model from learning the syntax and semantics meaning of primitives, NeSymReS (Luca et al., 2021) utilizes the pre-order traversal of binary tree to encode the expression. The constants in the expression are replaced with a placeholder “$c$” and optimized later after the skeleton is sampled according to the probability distribution outputted by the network. The pre-trained large-scale models have fast inference speed to obtain the predicted expression when confronted with new SR problems. The critical challenge of the SR task lies in two aspects: determining the skeleton and optimizing the constants. The above-stated large-scale pre-trained model could do well in predicting the skeleton, however, they normally use BFGS to optimize the constants which are sensitive to initial guesses. Therefore, to determine the constants accurately, based on the NeSymRe framework, Symformer (Vastl et al., 2022) and End2End (Kamienny et al., 2022) design two types of constant encoding strategies to teach the pre-trained model the exact value of constants. They perform with outstanding fitting accuracy than NeSymReS. Beyond the constant optimizing problem, Li et al (Li et al., 2022b) uses the Intra-class contrastive loss to train a better data encoder to tackle the ill-posed problem in SR. UDSR (Landajuela et al., 2022), SNR (Liu et al., 2023), and DGSR(Holt et al., 2023) merged multiple learning strategies to give the pre-trained model a revised mechanism. Moreover, the Monte Carlo Tree Search (Browne et al., 2012) is incorporated into the pre-trained model to achieve better performances.e.g.(Li et al., 2024). However, all the works view the SR task as a machine translation task that considers the input and output of the pre-trained model are the same data type. They ignore the fact that the data points are set data and the expression label is text. Thus, our work designs new architecture to solve the modality problem. SNIP(Meidani et al., 2023) employs a two-stage strategy to train a symbolic regression model. We pre-train an encoder using contrastive learning and then train a decoder separately. Since the two-stage training strategy is not conducive to feature fusion, SNIP also needs to introduce the LSO method to work well.

### v5

# MMSR: Symbolic Regression is a Multi-Modal Information Fusion Task

@@ Abstract
Mathematical formulas are the crystallization of human wisdom in exploring the laws of nature for thousands of years. Describing the complex laws of nature with a concise mathematical formula is a constant pursuit of scientists and a great challenge for artificial intelligence. This field is called symbolic regression (SR). Symbolic regression was originally formulated as a combinatorial optimization problem, and Genetic Programming (GP) and Reinforcement Learning algorithms were used to solve it. However, GP is sensitive to hyperparameters, and these two types of algorithms are inefficient. To solve this problem, researchers treat the mapping from data to expressions as a translation problem. And the corresponding large-scale pre-trained model is introduced. However, the data and expression skeletons do not have very clear word correspondences as the two languages do. Instead, they are more like two modalities (e.g., image and text). Therefore, in this paper, we proposed MMSR. The SR problem is solved as a pure multi-modal problem, and contrastive learning is also introduced in the training process for modal alignment to facilitate later modal feature fusion. It is worth noting that to better promote the modal feature fusion, we adopt the strategy of training contrastive learning loss and other losses at the same time, which only needs one-step training, instead of training contrastive learning loss first and then training other losses. Because our experiments prove training together can make the feature extraction module and feature fusion module wearing-in better. Experimental results show that compared with multiple large-scale pre-training baselines, MMSR achieves the most advanced results on multiple mainstream datasets including SRBench. Our code is open source at https://github.com/1716757342/MMSR.
keywords:
Symbolic Regression , Multi-modal , Information fusion , Contrastive learning , Modal alignment
journal: Nuclear Physics Bfn1fn1footnotetext: The first two authors contributed equally to this work.
[…]
[4]organization=School of Electronic, Electrical and Communication Engineering, University of Chinese Academy of Science, city=Beijing, postcode=101408, country=China
{graphicalabstract}
{highlights}
Propose a new multi-modal-based approach for symbolic regression.
Achieve an effective modality alignment mechanism utilizing contrastive learning.
Design a novel joint training strategy based on multi-loss fuction.
Our code is open source at https://github.com/1716757342/MMSR.
@@ 1 Introduction
Symbolic Regression (SR) aims to identify the explicit mathematical expression underlying the observed data. Formally, for a given dataset $D=\{X,y\}$, SR seeks a function $f$ that satisfies $y=f(X)$, where $X\in\mathbb{R}^{n\times d}$, $y\in\mathbb{R}^{n}$, $d,n$ are the dimension of variable and number of data points, respectively. $f$ is composed of several basic primitive operators such as +,,×,÷+,-$\times$,$\div$+ , - , × , ÷, etc. SR has been applied in many fields, for example, material sciences[1], physical law exploration[2][3][4], etc.
@@
In symbolic regression, expressions are usually represented as binary trees. Genetic Programming Symbolic Regression (GPSR)[5][6][7][8][9][10] has dominated in the past. GPSR uses the pre-order traversal of expression and acts as genes, and genetic operations such as mutation and crossover are used to produce descendants. Constants are randomly generated or optimized with the Particle Swarm Optimization (PSO)[11]. Although they have shown distinguished performances in several public benchmarks, they are sensitive to hyperparameters [12] and the complexity of the learned expression tends to be high with the epoch increased.
@@
With the huge success gained by deep learning in multiple domains [13], some works apply the deep learning technique in SR. Equation learner (EQL) [14] first substitutes the activation function of a fully connected feedforward neural network and refines the expression with Lasso. However, EQL is hard to tackle the division operation because it would generate exaggerated gradients in backpropagation. Therefore, Deep Symbolic Regression (DSR) [12] utilizes an RNN to emit the skeleton of expressions, constants are represented with a constant placeholder and optimized later by the numerical optimization algorithm. The risk-seeking policy gradients are used to update the RNN. To step further, Neural-Guided Genetic Programming (DSO, NGGP) [15] incorporates genetic programming into the search, RNN provides a better population for genetic programming algorithms. By doing so, DSO achieves remarkable results in SRBench [16], Since the efficiency of reinforcement learning is lower than that of numerical optimization, MetaSymNet[17] skillfully uses numerical optimization problems to solve the combinatorial optimization problem of symbolic regression.
Figure 1: Multi-model-based Symbolic Regression task versus multi-model-based Image Caption task. (a) Multi-model-based Symbolic regression task. In symbolic Regression task, an expression is expected to get from a set of input data [X,Y]. It can be considered as a multi-modal task, where the expression and the data are two different modalities. (b) Multi-model-based Image Caption task. The text modality is output to caption an image.
@@
The above-stated methods belong to the end-to-end model which trains a new model for each problem. This type of training strategy makes their search time long. Plus, the model can’t make full use of the search experiences from other searches. Therefore, the large-scale pre-trained model is proposed to improve the search speed. Normally, the large-scale pre-trained models [18][19][20] treat the symbolic regression problem as a language translation problem and train with enormous self-generated data which contains the pairwise of sampled data points and corresponding mathematical expressions skeletons. The neural network has an encoder-decoder structure, the encoder maps the input data points into data features and the decoder uses the feature to generate the pre-order traversal of expressions. The Beam Search [21] technique is applied to produce multiple candidate skeletons. Then, Broyden-Fletcher-Goldfarb-Shanno (BFGS) [22] is used to optimize the constants of candidate skeletons. Finally, the best-performed expression is selected as the final result.
@@
The large-scale pre-trained model gained fast search speed because once the model finished training, we only needed a single forward pass to get the expressions. However, they ignore the different data formats of the input and output of SR. That is, previous methods treat the SR as a machine translation task: translate the data points into corresponding expressions. However, it is clear that the input and output of SR are very different from the two languages in translation, and they do not have obvious word correspondence like languages. They are more like two different modalities (like image and text). Treating an SR question as a multi-modal task and cross-referenced this process with an image caption task is illustrated in Fig. 1. Therefore, it is perhaps more reasonable to think of SR as a multi-modal task.
@@
To tackle the problems of two different modalities, we use a modal alignment module in our model to make the encoded data feature close to the skeleton feature. We name our model as MMSR. Specifically, we use the SetTransformer [23] as a data encoder for it presents permutation invariance. And, decodes the skeleton of expression with a causal masking transformer decoder. Unlike standard decoder transformers, MMSR omits cross-attention in the first half of the decoder layers to encode unimodal skeleton representations, and cascades the rest of the decoder layers, cross-attending to the data encoder for multi-modal data-skeleton representations. We use contrastive learning [24] to align the modal features, so that the matching data features and skeleton features are as close as possible in the feature space, and the non-matching features are as far away from each other as possible. Additionally, we train the model with two extra losses. First, the cross-entropy loss is used for predicting the token of expression skeletons. Second, one of the limitations of pre-train models is that they focus on the supervised pretraining goals borrowed from symbol generation, i.e., they are trained solely with the token-level cross-entropy (CE) loss, which can result in equations that may exhibit high token-level similarities but are suboptimal with respect to equation-specific objectives such as fitting accuracy[25]. To solve this problem, we directly predict the numerical value of the constants and use MSE to measure the loss between the predicted constants and the true constants in the skeletons. The main contributions are summarized as follows:
  • 1.
    Propose a new multi-modal-based approach for symbolic regression. The idea of solving a symbolic regression problem via modal alignment and modal feature fusion significantly enhances the ability to accurately obtain concise expressions from data.
  • 2.
    Achieve an effective modality alignment mechanism utilizing contrastive learning. In this mechanism, deep-level feature representations are derived by contrasting positive and negative samples, and subsequently mapped them into a feature space that matches data and expressions. It promotes an enhanced understanding and differentiation of the differences within the data, thereby significantly improving the model’s comprehension capability regarding data and expressions.
  • 3.
    Design a novel joint training strategy based on multi-loss fuction. This strategy can adaptively achieve the optimal balance of expression skeleton, modal features, and data fitting accuracy under the collaborative constraints of cross-entropy loss ($\mathcal{L_{CE}}$), mean squared error (MSE) loss ( $\mathcal{L_{MSE}}$ ), and contrastive loss ( $\mathcal{L_{CON}}$ ), without over-relying on a single loss function.
@@ 2.1 Multimodal
Recently, models such as Contrastive Language-Image Pre-Training (CLIP) [26] and ALIGN [27] performed pre-train on noisy images and text pairs from the web using contrastive loss. Contrastive loss is one of the most effective feature learning methods[28][29][30][31]. They achieve remarkable performance on image-text retrieval tasks but lack the ability to model more complex interactions between image and text for other V+L tasks [32]. (e.g. VQA[33]). Subsequent investigations [34] [35] [36] have introduced encoder-decoder frameworks trained utilizing generative loss functions, which demonstrate robust performance across vision-language benchmarks. Concurrently, the visual encoders within these models maintain competitive accuracy in image classification tasks. Studies [37] [38] [39][40][41] have investigated the unification of image and text representations, which require multiple pretraining stages of unimodal and multi-modal modules to achieve high-performance levels. For example, ALBEF [38] employs a dual-encoder architecture that integrates contrastive loss with Masked Language Modeling (MLM) to enhance learning efficiency. CoCa[42] focuses on training an image-text foundation model from scratch in a single pretraining stage to unify these approaches. This is simpler and more efficient to train. BEITv3[43] regards images as a kind of language, and maps images into the language space through a mapping layer, and then sends them into the GPT large model together with the encoded text features. LLava[44] is an open-source multi-modal large model that uses images as a language. First, the big oracle model is frozen, and pre-trained, and the picture features are aligned with the text features to turn the picture into a ’language’ that the large language model(LLM)[45][46][47][48][49] can understand, and then concatenated with the text features and sent to the large language model.
@@ 2.2 Deep Symbolic Regression
Before the explosion of deep symbolic regression methods, Genetic Programming Symbolic Regression (GPSR) [50][51][52][53] was the main strategy for solving symbolic regression tasks. GPSR encodes the pre-order traversal of mathematical expression as the gene type, by randomly initializing a population of individuals, GPSR generates the candidate set of expressions. Then, it uses a fitness function to evaluate the fitting results, and genetic operations such as mutation and crossover are applied to produce new generations. The algorithm iterates until reaches the expected fitness value or the pre-defined maximum iteration epoch. Although GPSR could perform well in public benchmarks, it is reported to be sensitive to hyperparameters [12]. Furthermore, the GP-based method lacks an effective way to determine the constant value in the expression. RSRM[54] skillfully combines GP, Monte Carlo Tree Search (MCTS), and Double Q-learning. MSDB block is introduced to define new operators heuristically, which greatly improves the running efficiency of the algorithm. With the huge success achieved by deep learning in various domains, EQL revised the activation functions of fully connected neural networks to primitives (sin, cos, etc.) and obtained the skeleton and constants through Lasso [55] technique. However, the nontraditional activation functions cause troubles to the backpropagation process: the division operator would cause infinite gradients thus making the weights updating hard to continue. Moreover, the primitive sin, cos are periodical, and log, sqrt are not differentiable everywhere. DSR [12] takes another learning strategy: it learns the skeleton first and optimizes the constants later. DSR uses an RNN to emit the probability distribution of the pre-order traversal, and autoregressively sample expressions according to the distribution. For gradient information lost in the sampling process, reinforcement learning is applied to update the RNN. Specifically, the risk-seeking policy gradients are used. Furthermore, to utilize the strong ability to generate new performed well expressions of GPSR, DSO [15] incorporating the DSR and GPSR method to achieve state-of-the-art performance in SRBench. The RNN could provide a better population for GPSR to produce new descendants.
@@
DSO is a type of learning-from-scratch method, it needs to train a new model for every problem. This training strategy comes with a huge search time cost. Thus, many learning-with-experience methods are emerged. They usually employ a data encoder to receive the data points as inputs and output the expression encoding as the label to train a large-scale model. When a new problem needs to be handled, the learning-with-experience method just implements a forward propagation to obtain the expressions thus being time-effective. SymbolicGPT [18] treats the problem as a natural language problem, it uses symbolic-level encoding as the label of a set of data points. For example, the expression $2sin(x+y)$ is represented as text $[^{\prime}2^{\prime},^{\prime}s^{\prime},^{\prime}i^{\prime},^{\prime}n^{\prime},^{\prime}(^{\prime},^{\prime}x^{\prime},^{\prime}+^{\prime},^{\prime}y^{\prime},^{\prime})^{\prime}]$. This type of encoding forces the model to learn the syntax and semantics of the primitives thus increasing the learning difficulty. To free the model from learning the syntax and semantics meaning of primitives, NeSymReS [19] utilizes the pre-order traversal of binary tree to encode the expression. The constants in the expression are replaced with a placeholder “$c$” and optimized later after the skeleton is sampled according to the probability distribution outputted by the network. The pre-trained large-scale models have fast inference speed to obtain the predicted expression when confronted with new SR problems. The critical challenge of the SR task lies in two aspects: determining the skeleton and optimizing the constants. The above-stated large-scale pre-trained model could do well in predicting the skeleton, however, they normally use BFGS to optimize the constants which are sensitive to initial guesses. Therefore, to determine the constants accurately, based on the NeSymRe framework, Symformer [20] and End2End [56] design two types of constant encoding strategies to teach the pre-trained model the exact value of constants. They perform with outstanding fitting accuracy compared to NeSymReS.
Beyond the constant optimizing problem, Li et al [57] uses the Intra-class contrastive loss to train a better data encoder to tackle the ill-posed problem in SR. UDSR [58], SNR [59], and DGSR[60] merged multiple learning strategies to give the pre-trained model a revised mechanism. Moreover, the Monte Carlo Tree Search [61] is incorporated into the pre-trained model to achieve better performances.e.g.[62]. However, all the works view the SR task as a machine translation task that considers the input and output of the pre-trained model as the same data type. They ignore the fact that the data points are set data and the expression label is text. Thus, our work designs new architecture to solve the modality problem. SNIP[63] employs a two-stage strategy to train a symbolic regression model. SNIP pre-trains an encoder using contrastive learning and then trains a decoder. Since the two-stage training strategy is not conducive to feature fusion, SNIP also needs to introduce the LSO method to work well. TPSR[25] combines the large-scale pre-trained model and MCTS and uses the pre-trained model as a policy network to guide MCTS to search. The search efficiency of MCTS is improved.